kenzie126:

Simplify 3 square root of 10 end root plus 7 square root of 15 end root minus 6 square root of 10 end root minus 4 square root of 15 . A- 3 square root of 10 end root minus 3 square root of 15 B- 3 square root of 15 end root minus 3 square root of 10 C- 3 square root of 30 end root minus 3 square root of 20 D- 3 square root of 20 end root minus 3 square root of 30

supie:

Does it look like this-? \(\sqrt[3]{10}+\sqrt[7]{15}-\sqrt[6]{10}-\sqrt[4]{15}\)

kenzie126:

idk

supie:

It doesn't show you the actual problem? it just says "Simplify 3 square root of 10 end root plus 7 square root of 15 end root minus 6 square root of 10 end root minus 4 square root of 15 . "

kenzie126:

yeah

callmedaddy22:

omg thats dumb

kenzie126:

i know